The square surface shown has a side length of 3.9 mm. It is in a uniform electric field with a magnitude of 1925 N/C, which makes an angle of 44.0 degrees with a normal to the square surface as shown. Take that normal to define outward direction for the surface. Calculate the electric flux through the surface.Normal 350
Answer #1

Area of a square

A=s2=(3.9*10-3)2=1.521*10-5 m2

The electric flux through the surface is given by

\phi =EACos\theta =1925\times (1.521\times 10^{-5})Cos(180-44)

\phi =-0.02106 N-m^{2}/C